OpenStudy (aminehayyad):

The width of a rectangle is 2cm less than the length. The perimeter is 200cm. Find the length and width. Write an equation to express the dimensions of the rectangle.

OpenStudy (anonymous):

The formula we use is P=2(l+w)

OpenStudy (anonymous):

l=P 2﹣w=200/ 2﹣2=98cm

OpenStudy (aminehayyad):

but I thought only the width is different and it is 2 less, so where does the times 2 come to play?

OpenStudy (anonymous):